A rod AB of mass M and length L is lying on a horizontal frictionless surface. A particle of mass m travelling along the surface hits the end ‘A’ of the rod with a velocity V0 in a direction perpendicular to AB. The collision is elastic. After the collision the particle comes to rest.
(a) Find the ratio m/M.
(b) A point P on the rod is at rest immediately after collision. Find the distance AP.
(c) Find the linear speed of the point P after a time pL/3V0 after the collision.

Grade:

2 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Jay,

after collision let linear and angular velocity of centre of rod be V and w. Let moment of inertia of rod perpendicular to rod through centre of rod be I,

mv0=MV,(v0*L/2)m=Iw

(m/2)(v0)2=(I/2)w2+(M/2)V2, solve the three equations to obtain V,w and M.

b)(lAPl-(L/2))w=V

c)at t=(pi)L/3V0 AP would have rotated by w*t

(linear speed)2=(V+((lAPl-(L/2))sin(2700-wt)))2+((lAPl-(L/2))cos(2700-wt))2

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Jay !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR

Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

after collision let linear and angular velocity of centre of rod be V and w. Let moment of inertia of rod perpendicular to rod through centre of rod be I,
mv0=MV,(v0*L/2)m=Iw
(m/2)(v0)2=(I/2)w2+(M/2)V2, solve the three equations to obtain V,w and M.

b)(lAPl-(L/2))w=V
c)at t=(pi)L/3V0 AP would have rotated by w*t
(linear speed)2=(V+((lAPl-(L/2))sin(2700-wt)))2+((lAPl-(L/2))cos(2700-wt))2

Thanks and Regards